Two geekie physics students with a combined mass of 127 kg jump into their old car to run out for some late night pizza. The distance between the front and back axles of the car is 2.9 m. When they get in the car, the springs compress a total of 9.9 cm. On their way to the Pizza Hut, when they go over a bump, the car oscillates up and down with a period of 1.6 seconds. 1)What is mass of the car? mcar =
Answer #1

stiffness k = mg / x = 127kg * 9.8m/s² / 0.099m = 12.57 kN/m

angular frequency ω = 2π / 1.6s = 3.93 rad/s

from ω = √(k/m) we get
m = k/ω² = 12.57*10^3N/m / (3.93rad/s)² = 813.86 kg
so the mass of the car is
M = (813.86- 127) kg = 686.9 kg.....Answer.

Hope this helps !

Please thumbs up.